# SPECIAL RELATIVITY

### v9

@@ 11 A Preferred Inertial Frame and the Second Order Experiments
It should be note that all well known methods to test the special relativity are second order optical measurements. The cardinal example is given by the Michelson-Morley interference experiment. A close examination of all second order optical experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars. It is not possible to detect the orbital velocity of the earth using second order interference experiments. This is consequence of the geometry of space-time.
A number of other experiments were curried out, which all attempted to measure the velocity of the earth relative to the sun. They are not connected with light interference phenomena. We shall discuss experiments that compare the rates of different atomic clocks that move together relative to the earth-based frame. The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame $S$ in the accelerated frame $S_{n}$ by means of the time dilation effect for moving reference (atomic) clock. Until now, all test of time dilation with atomic clocks in the earth-based frame have involved circle path geometry. According to the special relativity, in any clock experiment which probe the influence of the orbital velocity one must use the linear (one-way) path geometry. These are experiments of decays of rapidly moving unstable particles. However, the modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
@@ 11.1 Second Order Optical Experiments and Special Relativity
The example of the time dilation effect is the resonance absorption of gamma rays in a Mossbauer rotating disk experiments.The measurements of the resonance absorption of gamma rays for the same relative velocity between the source and the observer give a blue shift when the source is at the center, and a red shift when it is at the tip. This experiment demonstrates that there is asymmetry in the transverse Doppler effect (1). These measurements also show that there is no a frequency shift when the source and observer are at the tip. The pulse from the source of one accelerated observer on the circle, which has red shift in the center of a circle due to time dilation, is perceived by diametrically opposite observer without frequency alteration due to his time dilation. This will take place despite the fact that each of the observer, having exchanged the pulse, is moving in the lab reference frame $S$ with velocity $v$, in opposing motion.
@@
At first site the transverse Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ by means of the Doppler red-blue shift of second order for moving sources. In particular, the question arises whether it is possible to ascertain the state of motion of the earth relative to the sun-based frame.
@@
Let us consider the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiments. Suppose that an absorber is placed in the center of a disk (which is at rest to the earth-based frame) and an emitter is placed on the edge of the disk. The emitter is moving round absorber along a circle with constant velocity $w$. In order to keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $w/c$ and $v/c$, where $v$ is the earth orbital velocity. The frequency of the emitter changes with the velocity $\vec{w}$ in a manner as it suggest itself from the transverse Doppler effect. Indeed, the inner frequency of the emitter depends on its velocity $\vec{w}$, so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$, then neglecting small terms $\omega(t_{n})=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]$. Since the physical clock is at rest in the earth-based frame, its reading of its proper time is $d\tau=(1-v^{2}/2)dt_{n}$. We expect therefore the radiation emitted with the constant frequency $\omega_{0}$ from the emitter to arrive in the absorber with a varying frequency $\omega(\tau)=\omega_{0}[1-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(\tau)/c^{2}]$. If the emitter is moving along a circle the frequency of radiation received disk center should fluctuate periodically. Then the measure of the absorption would be expected to vary with the position of the emitter relative to the disk center.
@@
In the actual experiment no effect on the absorption was found, when the disk was made to rotate. The negative result of the above experiment can be understood supposing that only a part of the whole frequency fluctuation is considered. A more careful analysis shows that one must take into account the change in frequency due to the so-called radial Doppler effect. We discuss a setup based on a perpendicular geometry and at first site the radial Doppler effect should be absent. We already know from our discussion in previous chapters that there is the aberration of light in the earth-based frame and the radial component of the emitter velocity is equal to $\vec{w}\cdot\vec{v}/c$. Thus we find that the emitted radiation falling on the disk center with periodically changing frequency $\delta\omega=\omega_{0}\vec{v}\cdot\vec{w}(\tau)/c^{2}$ will be exactly in resonance with the changing inner frequency. The two types of effects compensate each other and no observable effect remains.
@@
The Doppler red shift of second order for moving atoms is a major prediction of special relativity. The first experimental confirmation of the transverse Doppler effect appeared in 1938. In this (Ives and Stilwell) experiment, ionized hydrogen molecules were accelerated in the ”cathode tube” up to 28 keV. The frequencies of the light emitted parallel and antiparallel to the beam direction were measured by a spectrograph (at rest in the laboratory). Now let us to analyze this experiment. We discuss a setup based on a parallel geometry. The radiation from the moving atoms, giving the Doppler shifted lines, was observed together with the radiation from the resting atoms existing in the same working volume, and giving an anshifted line. In that way Ives and Stilwell replaced the difficult problem of the determination of the asymmetry of shift of the ”red” and blue shifted lines with respect to the unshifted line.
@@
At first site, the orbital motion of the earth relative to the sun can be observed through this type of experiment (2). Indeed, if the inner frequency of the atom depends on its velocity $\vec{w}$ so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}-\vec{w}|^{2}}$ and the frequency of radiation received spectrograph should depend on the orbital velocity. In rotating disk experiments the effect which occurred on the ground of the transverse Doppler effect is compensated by the other effect. In the case of perpendicular geometry we explained of the results of the experiments taking into account the aberration of light together with the radial (i.e. classical) Doppler effect. We know that in the collinear geometry the aberration of light is absent and one must take into account only the radial Doppler effect. If we analyze the geometry of the situation, we find that the observed frequency (i.e frequency of the radiation falling on the detector) is $\omega=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]/[1-w/(c-\vec{w}\cdot\vec{v}/w)]$, where (according to Langevin metric) $c^{*}=c-\vec{w}\cdot\vec{v}/w$ is the coordinate velocity of light in the earth-based frame. Thus we find just as in the case of the rotating disk experiment that also in measurements of the Doppler red shift of second order for moving atoms the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains.
@@
Usually, such a cancellation is found to stem from a deep underlying principle. From the series of negative results like that obtained by Michelson and Morley it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity $v$ remained unsuccessful. And we should accept the view that there exist a general law of nature which prevents us to determine $v$ by laboratory experiments. Nevertheless, in this case there does not appear to be any such profound implication. This is a specific feature of the optical experiments of second order. Indeed, we already know from our discussion in previous chapters that it is possible to detect the orbital velocity without looking at anything external to the earth using aberration of light phenomenon.
Now many people who learn special relativity in the usual way find this disturbing. The argument that aberration shift must be symmetrical runs some thing like this (3): It is always possible to chose such variables, in which metric of the accelerated frame will be diagonal. The validity of Maxwell’s laws in all inertial frames implies that light propagates with the same velocity $c$ in all inertial frames. It implies the elimination of the privilege frame, so meaning that the notion of absolute motion does not have physical content at all.
What is wrong with a principle that the metric supplies all information about the physics of the situation, as described in the given coordinates? We already discussed this subject in the Chapter 5. At first glance, after the diagonalization of the Langevin metric, we have the symmetry between inertial frames. Where does the asymmetry comes from? The electrodynamics equations are not supplies all information about the physics of the situation. To solve the electrodynamics equations it is necessary to determine the initial conditions. We already discussed how one can transform the absolute time coordinatization to Lorentz coordinatization.The time under the Einstein synchronization is readily obtained by introducing the time offset factor. This time shift has the effect of rotation of the plane of simultaneity. As a consequence of this, the plane wavefront rotates in the accelerated frame after metric diagonalization. Now the information about the direction of observer acceleration is recorded in the radiation wavefront orientation.
The electrodynamics equations are Lorentz covariant. This is a consequence of the (pseudo-Euclidean) geometry of space-time. However, from the mathematical viewpoint, the Lorentz covariance of electrodynamics equations is insufficient to guarantee the covariant solutions. In contrast, the accepted in previous literature incorrect assumption that covariant equations must have covariant solutions has always been considered obvious. We will discuss this subject further in the Chapter 12.
@@ 11.2 A Specific Feature of the Second Order Optical Experiments
We have seen that the principle of relativity is experimentally confirmed only in restricted sense. Presently we find that in the second order experiments the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains. And we may accept the view that there exist a specific feature of the second order experiments which prevents us to determine the velocity of the earth relative to the sun by laboratory experiments. Let us verify that this assertion is correct. We can summarize all the effects that we shall now discuss by remarking that they have to do with the interference effects of two light beams. The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern. The two beam originate in the same source. Thus, it is not possible to detect the orbital velocity using second order (interference) experiments. Phase is 4D invariant (i.e. it is simply number). It is independent on the chosen inertial frame and on the chosen coordinatization. This is a consequence of the geometry of space-time. In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well defined physical phenomena and only phase is meaningful quantity. This discussion clearly show that the special relativity calculations yields the observed null fringe shift in the earth-based frame.
@@
Until now we have considered only the case in which we have deal with interference pattern as in Michelson interferometer. Now we shall more general and study the case of the Doppler shift of second order for moving atoms. The frequencies of the emitted light were measured by a grating spectrograph. There is no specific, important physical difference between diffraction grating and interferometer. Grating spectrometer is also based on the interference of incident and reflected light waves.
@@
A more interesting example is the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiments. We shall try to understand the effect in a very simple case. A source is placed a large distance away from a thin plate of absorber. We inquire about the field at a large distance on the opposite side of the absorber. According to the electrodynamics, an electric field is the vector sum of the fields produced by the external source and the fields produced by each charges in the plane of absorber. We know that absorber consists of atoms which contains nuclear. When the electric field of the source acts on the nuclear it drives the nuclear charge. And moving charges generate a field - they constitute new radiators. This means that the two plane waves which interfere are propagate in the same direction. We have in this case interference of the incident and (coherently) forward scattered waves. Our model of the nuclear oscillator has some damping force. Because we put into take account of damping, the index of refraction is now a complex number and we have destructive interference. In this case we deal also with interference phenomenon.
11.3 Time Dilation Effect for Moving Reference Atomic Clock
The next thing that we have to discuss is the time dilation effect. We have already discussed this subject in Section 9.3. Here’s what we found there:
1. We have found that the slowly down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured. In other words, the time dilation effect is absolute, but not relative.
2. According to the asymmetry between (initial) inertial and accelerated frames, there is a remarkable prediction on the theory of special relativity. Namely, suppose that frame $S$ is at rest with respect to the fixed stars and frame $S_{n}$ is accelerated with respect to the inertial frame $S$ up to the speed $v$. Now two clocks symmetrically accelerated in the frame $S_{n}$ up to the the same speed $v$ but moving in opposite direction. However, there is a difference in time dilation of accelerated clocks in the frame $S_{n}$. It is clear that clocks have different accelerational histories with respect to the fixed stars.
We would now like to describe an apparent paradox. The argument that all inertial frames are equivalent runs something like this. We know that in order to keep a Lorentz coordinates in the accelerated frame we need to perform a metric diagonalization. Here the metric in both reference frames will be diagonal, and, according to textbooks, the coordinates in this frames should be related by the Lorentz transformation. At first site the problem is complete symmetrical and both frames are equivalent. However, there is a real difference between the accelerated inertial frame and the inertial frame without accelerational history. Indeed, we should expect the metric tensor to be changed. The difference between the inertial and accelerated frames is understandable. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the value $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the value $g_{00}=1,g_{01}=0$. We must conclude that when we are dealing with (Lorentz ) transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames.
The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ by means of the time dilation effect for moving reference (atomic) clock. At first site it may appear as if orbital motion of the earth could be observed by time dilation experiments. Although in principle this can be done it is necessary to analyze the question in greater detail.
The proper time interval may be identified with particle lifetime, atomic transition periods, nuclear half-lives, but the crucial point here is that this interval is completely determined by his movement in the initial inertial frame. In 1971 Hafele and Keating synchronized four atomic clocks with reference clock at the laboratory, and then flew them around the earth in commercial jets (4) . The clocks were flown first from east to west and then from west to east. When the clocks returned to the laboratory they were compared with the reference clock. The clocks flown in the plane ran slower than the clocks that remainder on the earth. The agreement between the observed and predicted discrepancies confirms the predictions of special relativity for the clock acceleration with respect to the earth. In the actual experiment no effect of earth orbital velocity on the clock dilation was found.
[…]
@@ 11.4 Wigner Rotation and the Trouton-Noble Experiment
Now we shall discuss other second order experiments. They are not connected with the propagation of light and can be understood in terms of general dynamical consideration. The most important of these experiments was the experiment of Trouton and Noble.
The Trouton-Noble experiment looks for the turning motion of a charged parallel plate capacitor suspended at rest in the frame of the earth in order to measure the earth’s motion relative to the sun (i.e. relative to the fixed stars). The main idea of the experiment can be described as follows.
@@
In the actual experiment a charged condense was suspended on an elastic string. The condenser was placed so that $\theta=\pi/4$, i.e. so that the line perpendicular to the surface of the condenser plates subtended an angle $\pi/4$ with the supposed direction of the orbital velocity of the earth, the direction of $\vec{v}$. This experiment led to negative result. i.e. the expected effect proportional to $v^{2}/c^{2}$ was not found to occur.
@@
The fact that capacitor in the earth-based frame produces no torque is in complete agreement with the special relativity and could be explained in a rather trivial way. We can obtain the electrodynamics equations in the earth-based frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. Here $\vec{v}$ is the velocity vector of the earth in the sun-based frame. Thus we must conclude that this magnetic field produces no torque in the earth-based frame where capacitor is at rest.
@@
In the experiment discussed we have still a puzzle. The existence of the magnetic field is responsible for the existence of the magnetic force and this force provides a torque on the charge in the sun-based frame. It is important to point out that experiment can be fully accounted for in terms of two inertial frames. We have the apparent paradox of different mechanical equations for force and torque governing the motion of a charged particle in different inertial frames; there is a torque and so a time rate of change of (3D) angular momentum in one inertial frame, but not in another relatively moving inertial frame. We will show that the real cause of the paradox is - the incorrect assumption that an earth-based observer and an sun-based observer have common 3-space. According to the special relativity, two observers with different trajectories have different 3-spaces. The standard presentation of the Trouton-Noble experiment is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving earth-based observer and $(x,y,z)$ axes of the sun-based observer are parallel. We demonstrate that only solution of the dynamics equations in covariant form (accounting for the special relativity in the transformation of coordinate axes) gives the explanation of Trouton-Noble experiment in a complete agreement with the special relativity.
@@
The key idea in resolution of the paradox arises in all usual explanation of the Trouton-Noble experiment is realization of the fact that in the space measurement of the earth-based observer there is mixing of space and time, as seen by the sun-based observer. While naively one should expect the angle between $\vec{v}$ and $\vec{r}_{n}$ in the earth-based frame to be the same angle between $\vec{v}$ and $\vec{r}$ in the sun-based frame it is not the case. The interpretation of the radius-vector rotation about the velocity vector is closely associated with the length contraction. Consider a rod directed along $x$-axis in the earth-based frame. After the Lorentz boost along the $x$-axis at velocity $v$ this rod will be contracted in the sun-based frame down to $l_{n}/\gamma$, where $l_{n}$ is the rod length in the earth-based frame. The relation $\gamma\tan\theta_{n}=\tan\theta$ gives the connection between the angle $\theta_{n}$ in the earth-based frame and the angle $\theta$ in the sun-based frame. Assuming $v^{2}/c^{2}\ll 1$, we find $\Delta\theta/\cos^{2}\theta=-[v^{2}/(2c^{2})]\tan\theta$, where $\Delta\theta=\theta_{n}-\theta$ is the rotation angle which is connected with the problem parameters by the relation $\Delta\theta=-[v^{2}/(2c^{2})]\sin\theta\cos\theta$. Thus we find that the moment of force produced by a pair of charges in the earth-based frame is reduced to $M_{n}=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$.

### v10

@@ 14.1 Second Order Optical Experiments and Special Relativity
An example of the time dilation effect is the resonance absorption of gamma rays in Mossbauer rotating disk experiments. The measurements of the resonance absorption of gamma rays for the same relative velocity between the source and the observer give a blue shift when the source is at the center and a red shift at the tip. This experiment demonstrates that there is an asymmetry in the transverse Doppler effect  [^54] The first experimental confirmation of the time dilation in the resonance absorption of gamma rays was reported by Champeney, Isaak, and Khan in 1965 [52].. These measurements also show no frequency shift when the source and observer are at the tip. The pulse from the source of one accelerated observer on the circle, which has a red shift in the center of a circle due to time dilation is perceived by a diametrically opposite observer without frequency alteration due to his time dilation. This will occur despite the fact that each of the observers, having exchanged the pulse, is moving in the lab reference frame $S$ with velocity $v$, in opposing motion.
@@
At first sight, the transverse Doppler effect problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ using the Doppler red-blue shift of second order for moving sources. In particular, the question arises whether it is possible to ascertain the state of motion of the earth relative to the sun-based frame.
@@
Let us consider the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiment. Suppose that an absorber is placed in the center of a disk (which is at rest to the earth-based frame) and an emitter is placed on the edge of the disk. The emitter is moving around the absorber along a circle with constant velocity $w$. To keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $w/c$ and $v/c$, where $v$ is the earth’s orbital velocity. The frequency of the emitter changes with the velocity $\vec{w}$ in a manner as it suggests itself from the transverse Doppler effect. Indeed, the inner frequency of the emitter depends on its velocity $\vec{w}$, so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$, then neglecting small terms $\omega(t_{n})=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(t_{n})/c^{2}]$. Since the physical clock is at rest in the earth-based frame, the reading of its proper time is $d\tau=(1-v^{2}/2)dt_{n}$. We expect therefore the radiation emitted with the constant frequency $\omega_{0}$ from the emitter to arrive in the absorber with a varying frequency $\omega(\tau)=\omega_{0}[1-w^{2}/(2c^{2})-\vec{v}\cdot\vec{w}(\tau)/c^{2}]$. If the emitter is moving along a circle the frequency of radiation received disk center should fluctuate periodically. Then the measure of the absorption would be expected to vary with the position of the emitter relative to the disk center.
@@
In the actual experiment, no effect on the absorption was found when the disk was made to rotate. The negative result of the above experiment can be understood supposing that only a part of the whole frequency fluctuation is considered. A more careful analysis shows that one must take into account the change in frequency due to the so-called radial Doppler effect. We discuss a setup based on a perpendicular geometry and at first sight, the radial Doppler effect should be absent. We already know from our discussion in previous sections that there is the aberration of light in the earth-based frame and the radial component of the emitter velocity is equal to $\vec{w}\cdot\vec{v}/c$. Thus we find that the emitted radiation falling on the disk center with periodically changing frequency $\delta\omega=\omega_{0}\vec{v}\cdot\vec{w}(\tau)/c^{2}$ will be exactly in resonance with the changing inner frequency. The two types of effects compensate each other and no observable effect remains.
@@
The Doppler red shift of second order for moving atoms is a major prediction of special relativity. The first experimental confirmation of the transverse Doppler effect appeared in 1938. In this (Ives and Stilwell) experiment, ionized hydrogen molecules were accelerated in the ”cathode tube” up to 28 keV. The frequencies of the light emitted parallel and antiparallel to the beam direction were measured by a spectrograph (at rest in the laboratory). Now let us analyze this experiment. We discuss a setup based on a parallel geometry. The radiation from the moving atoms, giving the Doppler shifted lines, was observed together with the radiation from the resting atoms existing in the same working volume, and giving an anshifted line. In that way, Ives and Stilwell replaced the difficult problem of the determination of the asymmetry of the shift of the ”red” and blue-shifted lines with respect to the unshifted line.
@@
At first sight, the orbital motion of the earth relative to the sun can be observed through this type of experiment  [^55] In Ives and Stilwell experiment [53] the effect which occurred on the ground of the orbital motion of the earth relative to the sun was within the experimental limitations. This situation pertains because of the high value of the emitter’s speed $w\gg v$ involved. Presently we give a general analysis of this type of experiment independently of the experimental accuracy.. Indeed, if the inner frequency of the atom depends on its velocity $\vec{w}$ so that (using Langevin metric in the earth-based frame) $\omega(t_{n})=\omega_{0}\sqrt{1-|\vec{v}+\vec{w}|^{2}}$ and the frequency of radiation received spectrograph should depend on the orbital velocity. In rotating disk experiments the effect which occurred on the ground of the transverse Doppler effect is compensated by the other effect. In the case of perpendicular geometry, we explained the results of the experiments taking into account the aberration of light together with the radial (i.e. classical) Doppler effect. We know that in the collinear geometry, the aberration of light is absent and one must take into account only the radial Doppler effect. If we analyze the geometry of the situation, we find that the observed frequency (i.e frequency of the radiation falling on the detector) is $\omega=\omega_{0}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})+\vec{v}\cdot\vec{w}(t_{n})/c^{2}]/[1-w/(c+\vec{w}\cdot\vec{v}/w)]$, where (according to Langevin metric) $c^{*}=c+\vec{w}\cdot\vec{v}/w$ is the coordinate velocity of light in the earth-based frame. Thus we find just as in the case of the rotating disk experiment that also in measurements of the Doppler red shift of second order for moving atoms the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains.
@@
Usually, such a cancellation is found to stem from a deep underlying principle. From the series of negative results like that obtained by Michelson and Morley, it seems reasonable to suppose that it is not accidental that various optical experiments trying to determine orbital velocity $v$ remained unsuccessful. And we should accept the view that there exists a general law of nature that prevents us from determining $v$ by laboratory experiments. Nevertheless, in this case, there does not appear to be any such profound implication. This is a specific feature of the optical experiments of second order. Indeed, we already know from our discussion in previous sections that it is possible to detect the orbital velocity without looking at anything external to the earth using aberration of light phenomenon.
@@ 14.2 A Specific Feature of the Second Order Optical Experiments
We have seen that the principle of relativity is experimentally confirmed only in a restricted sense. Presently we find that in the second order experiments, the two types of effects caused by the orbital motion of the earth compensate each other and no observable effect remains. We may accept the view that there exists a specific feature of the second-order experiments that prevents us from determining the velocity of the earth relative to the sun by laboratory experiments. Let us verify that this assertion is correct. We can summarize all the effects that we shall now discuss by remarking that they have to do with the interference effects of two light beams. The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern. The two beams originate from the same source. Thus, it is not possible to detect the orbital velocity using second-order (interference) experiments. Phase is 4D invariant. It is independent of the chosen inertial frame and the chosen coordinatization. This is a consequence of the geometry of space-time.
@@
This discussion clearly shows that the special relativity calculations yield the observed null fringe shift in the earth-based frame. The most famous of these interference experiments is one performed by Michelson and Morley. Now we shall study the case of the Doppler shift of second order for moving atoms. The frequencies of the emitted light were measured by a grating spectrograph. There is no specific, important physical difference between diffraction grating and interferometer.
@@
A more interesting example is the measurements of the resonance absorption of gamma rays in a Mossbauer rotating disk experiment. We shall try to understand the effect in a very simple case. A source is placed a large distance away from a thin plate of absorber. We inquire about the field at a large distance on the opposite side of the absorber. According to electrodynamics, an electric field is the vector sum of the fields produced by the external source and the fields produced by each charge in the plane of the absorber. We know that absorber consists of atoms that contain nuclear. When the electric field of the source acts on the nuclear it drives the nuclear charge. And moving charges generate a field - they constitute new radiators. This means that the two plane waves which interfere are propagating in the same direction. We have in this case interference of the incident and (coherently) forward scattered waves. Our model of the nuclear oscillator has some damping force. Because we put into take account damping, the index of refraction is now a complex number and we have destructive interference. In this case, we deal also with the interference phenomenon.
@@ 14.3 Wigner Rotation and the Trouton-Noble Experiment
Now we shall discuss other second-order experiments. They are not connected with the propagation of light and can be understood in terms of general dynamical consideration. The most important of these experiments was the experiment of Trouton and Noble [54]. The Trouton-Noble experiment looks for the turning motion of a charged parallel plate capacitor suspended at rest in the frame of the earth in order to measure the earth’s motion relative to the sun (i.e. relative to the fixed stars). The main idea of the experiment can be described as follows.
@@
In the actual experiment, a charged condense was suspended on an elastic string. The condenser was placed so that $\theta=\pi/4$, i.e. so that the line perpendicular to the surface of the condenser plates subtended an angle $\pi/4$ with the supposed direction of the orbital velocity of the earth, the direction of $\vec{v}$. This experiment led to negative results. i.e. the expected effect proportional to $v^{2}/c^{2}$ was not found to occur.
@@
The fact that capacitor in the earth-based frame produces no torque is in complete agreement with the special relativity and could be explained in a rather trivial way. We can obtain the electrodynamics equations in the earth-based frame using the Galilean transformation (with velocity $-v$) of Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. Here $\vec{v}$ is the velocity vector of the earth in the sun-based frame. Thus we must conclude that this magnetic field produces no torque in the earth-based frame where the capacitor is at rest.
@@
The existence of the magnetic field is responsible for the existence of the magnetic force and this force provides a torque on the charge in the sun-based frame. We have the apparent paradox of different mechanical equations for force and torque governing the motion of a charged particle in different inertial frames; there is a torque and so a time rate of change of (3D) angular momentum in one inertial frame, but not in another relatively moving inertial frame. We will show that the real cause of the paradox is - the incorrect assumption that an earth-based observer and a sun-based observer have a common 3-space. According to the special relativity, two observers with different trajectories have different 3-spaces. The standard presentation of the Trouton-Noble experiment is based on the hidden assumption that $(x_{n},y_{n},z_{n})$ axes of the moving earth-based observer and $(x,y,z)$ axes of the sun-based observer are parallel. We demonstrate that only the solution of the dynamics equations in covariant form (accounting for the special relativity in the transformation of coordinate axes) gives the explanation of the Trouton-Noble experiment in complete agreement with the special relativity.
@@
The key idea in the resolution of the paradox that arises in all usual explanations of the Trouton-Noble experiment is the realization of the fact that in the space measurement of the earth-based observer there is the mixing of space and time, as seen by the sun-based observer. While naively one should expect the angle between $\vec{v}$ and $\vec{r}_{n}$ in the earth-based frame to be the same angle between $\vec{v}$ and $\vec{r}$ in the sun-based frame it is not the case. The interpretation of the radius-vector rotation about the velocity vector is closely associated with the length contraction. Consider a rod directed along $x$-axis in the earth-based frame. After the Lorentz boost along the $x$-axis at velocity $v$ this rod will be contracted in the sun-based frame down to $l_{n}/\gamma$, where $l_{n}$ is the rod length in the earth-based frame. The relation $\gamma\tan\theta_{n}=\tan\theta$ gives the connection between the angle $\theta_{n}$ in the earth-based frame and the angle $\theta$ in the sun-based frame. Assuming $v^{2}/c^{2}\ll 1$, we find $\Delta\theta/\cos^{2}\theta=-[v^{2}/(2c^{2})]\tan\theta$, where $\Delta\theta=\theta_{n}-\theta$ is the rotation angle which is connected with the problem parameters by the relation $\Delta\theta=-[v^{2}/(2c^{2})]\sin\theta\cos\theta$. Thus we find that the moment of force produced by a pair of charges in the earth-based frame is reduced to $M_{n}=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$.
But we are still not finished! We must make a further correction to the moment of force in the earth-based frame. A more careful analysis shows that we must take into account the change in $v_{\perp}=v\sin\theta$ due to the Wigner rotation. As the force acts in the direction $\vec{r}$ there is a component of the charge acceleration perpendicular to the velocity $\vec{v}$. This acceleration component produces a rate of rotation of the proper (earth-base) frame axes in the sun-based frame. The rate at which the Wigner rotation occur is given $d\vec{\Phi}/dt=[v^{2}/(2c^{2})](\vec{v}/v^{2})\times d\vec{v}/dt$. The direction of the earth-based frame rotation in the sun-based frame is the same as the direction of the velocity rotation in the sun-based frame. We already have all the formulas that we need. For $d\theta/dt$ we get $d\theta/dt=[v^{2}/(2c^{2})]e^{2}\sin\theta/(mvr^{2})$, where $m$ is the mass of charge. So we have that the second correction to the moment of force is $mrdv_{\perp}/dt=v\cos\theta d\theta/dt=e^{2}v^{2}\sin\theta\cos\theta/(2c^{2}r)$. The two types of effects compensate the influence of the magnetic field in the sun-base frame and the moment of force produced by a pair of charges in the earth-based frame vanishes.
14.4 Time Dilation Effect for Moving Reference Atomic Clock
The next thing that we have to discuss is the time dilation effect. We have already discussed this subject in the Chapter 12. Here’s what we found there:
1. We have found that the slow down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of the reference system, in which this effect is measured. In other words, the time dilation effect is absolute, but not relative.
2. According to the asymmetry between (initial) inertial and accelerated frames, there is a remarkable prediction in the theory of special relativity. Namely, suppose that frame $S$ is at rest with respect to the fixed stars and frame $S_{n}$ is accelerated with respect to the inertial frame $S$ up to the speed $v$. Now two clocks symmetrically accelerated in the frame $S_{n}$ up to the the same speed $v$ but moving in opposite direction. However, there is a difference in time dilation of accelerated clocks in the frame $S_{n}$. It is clear that clocks have different accelerational histories with respect to the fixed stars.
The question arises whether it is possible to determine experimentally the state of motion of the inertial frame $S$ in the accelerated frame $S_{n}$ using the time dilation effect for moving reference (atomic) clock. At first sight, it may appear as if orbital motion of the earth could be observed by time dilation experiments. Although in principle this can be done it is necessary to analyze the question in greater detail.
The proper time interval may be identified with particle lifetime, atomic transition periods, and nuclear half-lives. Still, the crucial point here is that this interval is completely determined by clock movement in the initial inertial frame. In 1971 Hafele and Keating synchronized four atomic clocks with reference clock at the laboratory and then flew them around the earth in commercial jets [^56] The first experimental confirmation of the time dilation predicted in the circle path geometry with atomic clocks and aircraft was reported by Hafele and Keating in 1971 [55]. A very interesting result has been obtained by the GPS team. The first example of the different inertial frames having the same accelerational histories is the global positioning system (GPS). The number of GPS satellites is around 80. The velocity of the GPS satellites with regard to the earth is 4 km/s. The experimental data of the GPS shows that the clocks in the satellites tick off time more slowly (7.1 $\mu$s every day) by the velocity. Asymmetry in clock run appears between the earth and the satellites. There are no asymmetries among the GPS satellites. That is, times are equal in every GPS satellite [56]. We see that the presence of some non-zero relative velocity cannot be a cause for the effect of time dilation. In this experiment, no effect of earth orbital velocity on the clock dilation was found. The negative result of the GPS experiment can be understood by remembering that in this situation the clocks also are moving around the earth (i.e. around the reference clock) along the circular path and no observable effect of orbital velocity remains.. The clocks were flown first from east to west and then from west to east. When the clocks returned to the laboratory they were compared with the reference clock. The clocks flown in the plane ran slower than the clocks that remained on the earth. The agreement between the observed and predicted discrepancies confirms the predictions of special relativity for the clock acceleration with respect to the earth. In the actual experiment, no effect of earth’s orbital velocity on the clock dilation was found.